OpenStudy (anonymous):

Solve the system: 3x^2 - 6y^2 = 9 and x^2 + 5y^2 = 53

sam (.sam.):

3x^2 - 6y^2 = 9 x^2 + 5y^2 = 53 Multiply 3 ---------------------------- Elimination 3x^2 - 6y^2 = 9 3x^2 + 15y^2 = 159 ============= -21y^2=-150 y^2=50/7 --------------- \[~~~~~~~~~~~~~~~~~~ y= \pm \sqrt{\frac{50}{7}}\] Substitute this 'y' values to 3x^2 - 6y^2 = 9 and find the other 2 values of x